How did you solve 0.9939 g/mL and how you come up with that formula and how you get your final answer? I need some clarification ^_^. Thank you so much
1
Expert's answer
2020-09-09T04:35:52-0400

p=m/V = 0.9939 g / 1 ml = 0.9939 g/ml.